Capacitor values can be assumed to be large enough in the small-signal analysis frequency band. (Hint: Consider separating the bias and small-signal signals.) 1. Find the bias current flowing into the transistor (λ=0) 2. Find the small-signal voltage gain Vout/Vin value. (λ=0.1V^(-1))

Vpp=20V T VBias=3V Rp=10k Rg=90k 2 Vout VIN - HE Rsig=10812 B=1mA/V2 Vth=1V Rs=50012
Answer #1


Voorov VBISH=3V P-10 kr &&B=90k Vinth VB. Bimal? VH - ۱۷ Vic Rs=suon D a=0 Assume Transistor is in saturation region. 2 1 = (

so, Ip = (m) (3-V5-0 į (im) (2-vs) Ź Cim) (4+1-4V5) clearly, I = Is so = Is - Vs-o 500 from and Vs (im) (44 V7-405) llo? Tooo

So, Tof Vs 500 ID 0-535898 500 Ip = 0.0010716 II = 16 of 16 MA (2) To rol- 20 (0.1) (1.0916m) ro= 9331.84 gm = 2 IP Vers VGH

so clearly, No e fom Vgs) (roll Ro) Vgs = van Vo --50 (roll) Vigis Vo. goo Cool120) Vin Vo Vin :-(1.4637m) (9331.84 X 10k) 93